# A Multi-Fidelity Control Variate Approach for Policy Gradient Estimation

### v3

@@ 6.2 Empirical Variance Reduction
Figure 2: Variance of policy gradient estimates for MFPG versus variants of the High-Fidelity Only baseline on Hopper-v3 (top), and the percentage of MFPG variance relative to single-fidelity counterparts of the same batch size (bottom). In high-fidelity data-scarce regimes with mild dynamics gaps, MFPG generally exhibits far lower PG variance than the single-fidelity counterparts.
@@
We use a Hopper task with $1.2\times$ friction as an example. Figure˜2 (top) shows the variance of PG estimates from MFPG versus a few variants of the High-Fidelity Only baseline (more specifically, we plot the variances of the scalar quantities $Z^{\pi_{\theta}}$ and $X_{\tau^{h}}^{\pi_{\theta}}$ before differentiation with respect to policy parameters). The plotted High-Fidelity Only variants include a baseline that has access to the same limited number of high-fidelity samples per policy update as MFPG (i.e., the same batch size), baselines with more high-fidelity samples (a batch size of $1000$ and $3000$), and variants of these High-Fidelity Only baselines that also use state-value function subtraction in an effort to variance reduction.
@@
6.3 Large-Scale Comparison Across Dynamics Variation Types and Levels
Figure 3: Final evaluation returns of MFPG versus baselines in high-fidelity data-scarce regimes. On Hopper-v3 and Walker2d-v3, under mild–moderate multi-fidelity dynamics gaps, MFPG reliably improves median performance over the High-Fidelity Only baseline and matches that of leading multi-fidelity methods, despite its simplicity and minimal tuning overhead. Under large dynamics gaps, all multi-fidelity approaches converge toward the High-Fidelity Only baseline performance. On the more challenging HalfCheetah-v3, training stability decreases; here we additionally report the AUC metric (Fig.˜4) to capture accumulated performance, as it may diverge from final return trends. In this task, MFPG ranks among the best-performing methods overall and demonstrates the strongest robustness (cf. Fig.˜4).
Figure 4: AUC of MFPG versus baselines in high-fidelity data-scarce regimes on HalfCheetah-v3 with friction variations. Under an extreme dynamics gap (5× friction), we additionally report evaluation return curves for all methods—MFPG is the only multi-fidelity approach that remains comparable to High-Fidelity Only in accumulated performance (AUC), while all multi-fidelity methods suffer severe degradation in terms of final return (cf., Fig.˜4).
This section reports the large-scale evaluation of MFPG against baseline methods under the task settings described in Section˜6.1 and supports our key insight I2. For clarity, we present in Figs.˜4 and 4 a representative subset of results—specifically, one dynamics variation type with multiple variation levels for each task. The complete results across all 39 task settings are provided in Appendix˜C.
Fig.˜4 presents the final evaluation return of all methods on 15 representative task settings (out of 39). Each row corresponds to one task, with the high-fidelity environment varying across multiple dynamics levels. For the simpler Hopper and Walker2d tasks, training is relatively stable, and the final evaluation return and AUC metrics are consistent; thus, we report only the final evaluation return. In contrast, the HalfCheetah task is more challenging, with training exhibiting greater fluctuations that occasionally reduce reward performance (baselines fluctuate far more substantially than MFPG). To capture accumulated learning performance more comprehensively in these settings, we report both the final return (Fig.˜4, row 3) and the AUC metric (Fig.˜4).
Hopper and Walker2d tasks. At a high level, when the multi-fidelity dynamics gap is mild—e.g., rows 1–2 with $0.8\times$ and $1.2\times$ gravity or friction in Fig.˜4—the median of MFPG improves reliably over High-Fidelity Only, narrowing the gap to the median of the oracle baseline (with privileged access to $15\times$ more high-fidelity data). As the dynamics gap grows but remains moderate (e.g., $0.5\times$ and $2.0\times$ gravity or friction in rows 1–2), where cross-fidelity correlation weakens, the median of MFPG still generally improves upon High-Fidelity Only, albeit with smaller margins. In most of these settings, the medians of other baselines also improve upon High-Fidelity Only and are competitive with one another. Notably, MFPG generally matches the performance of leading multi-fidelity methods despite its algorithmic simplicity and minimal tuning overhead. In settings with large dynamics gaps, i.e., $5\times$ shifts, the performance of all evaluated approaches becomes generally comparable to High-Fidelity Only.
Another interesting observation is that the performance of Low-Fidelity Only (100×) does not consistently correlate with the dynamics shift level. For instance, in Hopper with 2× gravity, Low-Fidelity Only (100×) outperforms High-Fidelity Only in median, whereas at 0.5× gravity it yields a lower median—highlighting the inherent difficulty of predicting when zero-shot transfer will be effective in practice.
@@ 6.3 Large-Scale Comparison Across Dynamics Variation Types and Levels
A similar trend in PAR performance is observable under other types of dynamics variation of HalfCheetah, cf., gravity and kinematic shifts in Figs.˜12 and 13 of Appendix˜C. In HalfCheetah with kinematic variation, PAR’s median performance drops sharply below that of MFPG at the medium level and collapses entirely at the hard level on both metrics, whereas MFPG closely tracks the median of the High-Fidelity Only baseline throughout.
Robustness. Across all 39 task settings, the numbers of performance collapses (in final return / AUC) are as follows: MFPG (2 / 0), PAR (4 / 2), Low-Fidelity Only (100×) (11 / 9), and DARC (12 / 10). Among all methods, MFPG exhibits the strongest robustness.
MFPG performance vs. correlation. We examine one task in detail to illustrate the qualitative behavior of MFPG under varying dynamics gaps. Figure˜5 presents evaluation-return curves of MFPG and High-Fidelity Only in the Hopper task with three levels of gravity variation (0.8×, 2.0×, and 5.0×), along with the estimated Pearson correlation coefficients between high- and low-fidelity policy gradient losses throughout training. In this task, the correlation between fidelities generally decreases as the dynamics shift becomes larger. Correspondingly, the median performance gains of MFPG over High-Fidelity Only are greater when the correlation is higher (smaller gap). In the extreme case of 5× gravity, the correlation drops sharply, and MFPG performance nearly coincides with that of High-Fidelity Only.
@@
Figure 5: Evaluation return curves of MFPG versus the High-Fidelity Only baseline on Hopper-v3, corresponding to Fig.˜4, under mild, moderate, and large gravity variations (left to right: 0.8×, 2.0×, 5.0×), together with estimated Pearson correlation coefficients between high- and low-fidelity policy gradient losses. MFPG performance increases as the cross-fidelity correlation strengthens.
@@ 6.4 Low-Fidelity Environment with a Misspecified Reward
This experiment is designed to support our key insight I3, that when high-fidelity samples are scarce, MFPG can remain effective even in the presence of reward misspecification in the low-fidelity samples.
@@
To demonstrate this point, we examine a situation in which the reward function in the low-fidelity environment is the negative of that from the high-fidelity environment. Figure˜7 reports the performance of MFPG in comparison to the High-Fidelity Only and Low-Fidelity Only baselines. All approaches in this example use the plain REINFORCE algorithm without state-value subtraction. The core idea extends to the case with value-function subtraction, except that separate value functions must be learned for the multi-fidelity environments because of the misspecified reward.
Figure 6: MFPG, High-Fidelity Only, and Low-Fidelity Only on Hopper-v3, where the low-fidelity environment is misspecified with a negated reward. Although low-fidelity data alone is ineffective for training, MFPG successfully leverages both fidelities to achieve substantial improvements over High-Fidelity Only in high-fidelity data-scarce regimes.
Figure 7: Final return (top) and AUC (bottom) of MFPG versus High-Fidelity Only on Hopper-v3 with 1.2× friction, across varying high-fidelity batch sizes (i.e., samples per update). MFPG yields substantial median performance gains over High-Fidelity Only in high-fidelity data-scarce regimes (batch size 100). When high-fidelity data is sufficient (e.g., batch size 2000), augmenting with abundant low-fidelity data provides no further benefit.
@@ 6.5 Additional Experiments
This section presents additional experiments that provide a more detailed examination of different aspects of the MFPG framework.
@@
MFPG performance with different amounts of high-fidelity samples. Our main evaluation in Section˜6.3 focuses on the high-fidelity data-scarce regime, which reflects many real-world scenarios where collecting large quantities of high-fidelity samples for each policy update is prohibitively expensive (e.g., in robotics or molecule design). In this section, we study how MFPG compares to the High-Fidelity Only baseline as the number of high-fidelity samples per policy update varies.
Figure˜7 reports final return (top) and AUC (bottom) in the Hopper task with 1.2× friction, trained for 1 million steps using MFPG with 100× low-fidelity data and High-Fidelity Only across different batch sizes (i.e., high-fidelity samples per update). Note that with a fixed total number of high-fidelity environment steps, the number of gradient updates varies across batch sizes; consequently, the reported metrics do not necessarily increase monotonically with larger batch sizes. In the data-scarce regime (batch size 100), the medians of MFPG are substantially greater than those of High-Fidelity Only. The margin of improvement decreases as more high-fidelity samples become available (e.g., batch size 200), and when high-fidelity samples are abundant (e.g., batch size 2000), MFPG closely matches High-Fidelity Only. These results highlight that MFPG is most effective in high-fidelity data-scarce settings.
Figure 8: Final return (top) and AUC (bottom) of MFPG with increasing amounts of low-fidelity data (light to dark orange) versus High-Fidelity Only (blue) on Hopper-v3 with 0.5× and 1.2× friction. Overall, incorporating additional low-fidelity samples yields greater benefits in mild-gap settings.
Figure 9: Final return (top) and AUC (bottom) of MFPG with varying exponential moving-average weights ($\eta_{\text{ma}}$ in Eq.˜2) versus High-Fidelity Only on Hopper-v3 with 0.5× and 1.2× friction. MFPG remains robust to the choice of $\eta_{\text{ma}}$ with medians consistently exceeding those of High-Fidelity Only.
@@
MFPG performance with different amounts of low-fidelity samples. We assume throughout this work that low-fidelity samples are cheap to collect, and all methods are given access to an abundant budget of low-fidelity data. Consequently, optimizing the efficiency of low-fidelity sample usage (Cutler et al., 2015) is not the focus of this work. Instead, this section qualitatively examines how the performance of MFPG varies with different amounts of low-fidelity data.
@@
In principle, the performance of MFPG is robust with respect to the amount of low-fidelity samples, so long as the low-fidelity sample amount is sufficient to estimate the low-fidelity sample mean $\hat{\mu}^{l}$. Figure˜9 reports the final return (top) and AUC (bottom) of MFPG with increasing amounts of low-fidelity samples (indicated from light to dark orange) for estimating low-fidelity sample mean ($\hat{\mu}^{l}$ in Fig.˜1), compared with High-Fidelity Only in Hopper with 0.5× and 1.2× friction. At a high level, when the dynamics gap is mild (e.g., 1.2× friction), increasing the number of low-fidelity samples tends to yield higher median performance, especially in terms of AUC. However, under larger dynamics gaps, no clear monotonic trend is observed. Overall, additional low-fidelity samples are most beneficial in mild-gap settings.
@@
Sensitivity to exponential moving-average weight. We further examine the sensitivity of MFPG to the choice of the exponential moving-average weight $\eta_{\text{ma}}$ in Eq.˜2. Figure˜9 reports the final return and AUC of MFPG with different $\eta_{\text{ma}}$ values in the Hopper task under 0.5× and 1.2× friction, alongside the High-Fidelity Only baseline. Overall, MFPG is relatively robust to the choice of $\eta_{\text{ma}}$, i.e., the medians do not fall below those of High-Fidelity Only baseline, and its median performance improves more substantially in settings with smaller dynamics gaps—consistent with the key insights from our baseline comparison in Section˜6.3.
Reparameterization trick. We examine the effect of the reparameterization trick described in Section˜4 on our MFPG method. Figure˜11 shows that when the dynamics gap is mild (cf., the 1.2× friction setting in Hopper, right panel), reparameterization substantially accelerates learning; in contrast, when the gap is larger (cf., the 0.5× friction setting, left panel), it offers much weaker benefit. This pattern is intuitive: when low-fidelity dynamics are close to those of the high-fidelity environment, stochasticity in action selection dominates; reparameterization controls this source of randomness and produces highly correlated rollouts. In the extreme case where the low-fidelity dynamics are perfect (and deterministic), reparameterization makes the low- and high-fidelity rollouts identical, so that $X^{\pi_{\theta}}_{\tau^{h}}$ and $X^{\pi{\theta}}_{\tau^{l}}$ cancel out. The policy gradient estimate then reduces to $\mu^{l}$, which is estimated with abundant (perfect) low-fidelity samples. By contrast, when the dynamics gap is large, discrepancies in transition dynamics dominate, and even with reparameterization, divergence between high- and low-fidelity rollouts persists, resulting in weaker correlation.
Figure 10: MFPG with and without the reparameterization trick versus High-Fidelity Only on Hopper-v3 with 0.5× (left) and 1.2× (right) friction. The reparameterization trick proves especially critical in mild-gap settings (e.g., 1.2× friction).
Figure 11: MFPG with and without periodic reconstraining of low-fidelity trajectories versus High-Fidelity Only on Hopper-v3 with 0.5× (left) and 1.2× (right) friction. Reconstraining can enhance MFPG performance in some settings (cf. right) but degrade it in others (cf. left). Accordingly, we do not employ this trick in our evaluations of MFPG.
Variance-bias trade-off. In practice, we observed that periodically reconstraining (the correlated) low-fidelity rollouts back to their high-fidelity counterparts can strengthen correlation and improve performance (e.g., the 1.2× friction setting in Fig.˜11, right panel). However, this procedure may also introduce bias into the policy gradient estimator, leading to degraded performance in certain cases (e.g., the 0.5× friction setting in Fig.˜11, left panel). These findings highlight a practical variance–bias trade-off. We do not adopt this reconstraining trick in our experiments. Developing systematic, unbiased mechanisms for drawing even more highly correlated samples across environments remains an important direction for future work.

### v4

@@ 6.2 Empirical Variance Reduction
We use a Hopper task with $1.2\times$ friction as an example. Figure˜2 (top) shows the variance of PG estimates from MFPG versus a few variants of the High-Fidelity Only baseline (more specifically, we plot the variances of the scalar quantities $Z^{\pi_{\theta}}$ and $X_{\tau^{h}}^{\pi_{\theta}}$ before differentiation with respect to policy parameters). The plotted High-Fidelity Only variants include a baseline that has access to the same limited number of high-fidelity samples per policy update as MFPG (i.e., the same batch size), baselines with more high-fidelity samples (a batch size of $1000$ and $3000$), and variants of these High-Fidelity Only baselines that also use state-value function subtraction in an effort to reduce variance.
@@
Figure 2: Variance of policy gradient estimates for MFPG versus variants of the High-Fidelity Only baseline on Hopper-v3 (top), and the percentage of MFPG variance relative to single-fidelity counterparts of the same batch size (bottom). Solid lines denote the median across seeds, and shaded regions indicate the corresponding maximum and minimum values. In high-fidelity data-scarce regimes with mild dynamics gaps, MFPG generally exhibits far lower PG variance than the single-fidelity counterparts.
@@
6.3 Baseline Comparison Across Dynamics Variation Types and Levels
This section reports the evaluation of MFPG against baseline methods under the task settings described in Section˜6.1 and supports our key insight I2.
Fig.˜3 presents the mean difference in final evaluation return of each method relative to the High-Fidelity Only baseline, measured as the difference in mean final evaluation return $\Delta=\text{mean}_{\text{method}}-\text{mean}_{\text{High-Fidelity Only}}$ (so 0 corresponds to High-Fidelity Only), on 15 representative task settings, with error bars indicating two-sided $95\%$ bootstrap confidence intervals for $\Delta$. Each row corresponds to one task, with the high-fidelity environment varying across multiple dynamics levels. For the simpler Hopper-gravity and Walker2d-friction tasks, abundant low-fidelity data tend to be neutral or beneficial and training is relatively stable, so the final evaluation return and AUC metrics are generally consistent; thus, we report only the final evaluation return. In contrast, the HalfCheetah-friction task is more challenging and low-fidelity data tend to be harmful: training exhibits greater fluctuations that occasionally reduce reward performance (baselines fluctuate far more substantially than MFPG). To capture accumulated learning performance more comprehensively in these settings, we report both the final return (Fig.˜3, row 3) and the AUC metric (Fig.˜4). We compute and visualize the AUC results analogously, i.e., as differences relative to High-Fidelity Only with two-sided $95\%$ bootstrap confidence intervals.
@@ 6.3 Baseline Comparison Across Dynamics Variation Types and Levels
Figure 4: Mean difference in AUC for each method in HalfCheetah–friction relative to High-Fidelity Only, computed analogously to Fig.˜3 (0 corresponds to High-Fidelity Only). Bars show the bootstrap point estimate of the mean difference, and error bars denote two-sided $95\%$ bootstrap confidence intervals. When low-fidelity data become harmful, MFPG exhibits the strongest robustness to performance degradation among the evaluated methods.
Hopper and Walker2d tasks. As shown in Fig.˜3 (rows 1-2), in these two tasks, abundant low-fidelity data generally tend to be neutral or beneficial: across all settings, the confidence intervals of the Low-Fidelity Only baseline are never strictly below 0. MFPG is the only evaluated method that reliably and consistently achieves statistically significant improvements when the dynamics gap is mild to moderate (0.5× to 2.0×)—i.e., with error bars strictly above 0 in all 8/8 such scenarios—by leveraging low-fidelity data and cross-fidelity correlation to complement the limited high-fidelity data.
No other off-dynamics RL or low-fidelity-only baseline achieves this level of consistency. For instance, in Hopper-v3 with a 0.5× gravity shift, the Low-Fidelity Only baseline shows neutral performance relative to High-Fidelity Only. Consequently, the evaluated off-dynamics RL baselines fail to significantly improve over High-Fidelity Only, whereas MFPG effectively combines high- and low-fidelity data, leveraging cross-fidelity correlations to achieve a statistically significant performance gain over the High-Fidelity Only baseline.
When low-fidelity data tend to be strongly beneficial, e.g., in Hopper-v3 with a 1.2× gravity shift and Walker2d-v3 with 0.8×–2.0× friction shifts, the evaluated off-dynamics RL baselines aggressively exploit the abundant low-fidelity data to boost performance, achieving significantly higher mean returns than High-Fidelity Only. By contrast, MFPG behaves more cautiously, providing reliable and consistent—yet still statistically significant—improvements over High-Fidelity Only.
We emphasize that this behavior of MFPG is by design: MFPG uses low-fidelity data solely as a variance-reduction tool, deliberately trading potentially aggressive (but risky) gains for reliability. This design choice underlies the consistency observed here and the strongest robustness that MFPG exhibits relative to the other evaluated off-dynamics RL methods in the HalfCheetah-v3 task presented below.
In extreme-shift scenarios, i.e., 5.0× dynamics shifts, cross-fidelity correlation drops sharply and MFPG performance becomes effectively neutral relative to High-Fidelity Only.
Another interesting observation is that, even when it improves over High-Fidelity Only, the Low-Fidelity Only baseline tends to exhibit high cross-seed variance, as reflected by its long error bars in Hopper-v3 and Walker2d-v3 with 0.8×–2.0× dynamics shifts in the box plots shown in Figs.˜14 and 15. Moreover, whether Low-Fidelity Only significantly improves over High-Fidelity Only does not correlate monotonically with the dynamics gap; for example, low-fidelity data appear more beneficial in Hopper-v3 with a 2.0× gravity shift than with a 0.5× shift. This highlights both the inherent difficulty of predicting when zero-shot transfer will be effective in practice and the importance of having reliable multi-fidelity approaches.
HalfCheetah task. As shown in Fig.˜3 (row 3) and Fig.˜4, in this task, low-fidelity data tend to be harmful: DARC and Low-Fidelity Only consistently and significantly degrade mean final return performance over High-Fidelity Only across mild-to-moderate gaps (0.5× to 2.0×), whereas PAR still manages to improve significantly over High-Fidelity Only and achieves the highest performance among the evaluated methods in these scenarios. However, PAR’s performance gains come at the cost of computationally expensive, per-task hyperparameter tuning. We empirically observed that PAR is highly sensitive to the reward-augmentation weight $\beta$; see Fig.˜18 in Appendix˜F. Even with such per-task tuning, PAR’s performance often exhibits large cross-seed variance with low minima (cf. 0.5× friction in Fig.˜15 and several additional cases in Figs.˜16 and 17).
By contrast, MFPG does not require such expensive, careful tuning; thanks to its algorithmic simplicity, we only tuned the exponential moving-average weight $\eta_{\text{ma}}$ in Eq.˜3 during our experiments. Across mild-to-moderate gaps, MFPG is robust to harmful low-fidelity data and remains effectively neutral with respect to High-Fidelity Only.
At an extreme dynamics shift, i.e., 5.0× friction, the performance of DARC, PAR, Low-Fidelity Only all drops sharply, with the confidence intervals for both mean final return and AUC far below 0. In this scenario, we observe that the mean return curve of MFPG (cf. Fig.˜4) initially closely tracks that of High-Fidelity Only before exhibiting some instability, resulting in a lower mean final return but relatively strong accumulated performance (i.e., AUC). By contrast, all other baselines struggle to achieve meaningful performance throughout training, suggesting that MFPG is the most robust among the evaluated off-dynamics RL and low-fidelity-only methods in this setting.
The observed instability of MFPG may stem from noise in the estimated control-variate coefficient $\hat{c}^{*}$, and designing more reliable estimation schemes constitutes an interesting direction for future work.
[…]
Taken together, these results support our key insight I2 regarding the reliability and robustness of MFPG.
[…]
@@
Figure 5: Evaluation return curves of MFPG versus the High-Fidelity Only baseline on Hopper-v3, corresponding to Fig.˜3, under mild, moderate, and large gravity variations (left to right: 0.8×, 2.0×, 5.0×), together with estimated Pearson correlation coefficients between high- and low-fidelity policy gradient losses. Solid lines indicate bootstrap point estimates (means), and shaded regions denote two-sided 95% bootstrap confidence intervals. Consistent with the correlation estimates, the gap between MFPG’s mean return and that of High-Fidelity Only is more pronounced when cross-fidelity correlation is stronger.
@@ 6.4 Low-Fidelity Environment with a Misspecified Reward
This experiment is designed to support our key insight I3, that when high-fidelity samples are scarce, MFPG can remain effective even in the presence of reward misspecification in the low-fidelity samples, e.g., when the low-fidelity reward is the negative (anti-correlated) version of the high-fidelity reward.
@@
To demonstrate this point, we examine a situation in which the reward function in the low-fidelity environment is the negative of that from the high-fidelity environment. Figure˜7 reports the evaluation return of MFPG in comparison to the High-Fidelity Only and Low-Fidelity Only baselines. Solid lines indicate bootstrap point estimates (means), and shaded regions denote two-sided 95% bootstrap confidence intervals. All approaches in this example use the plain REINFORCE algorithm without state-value subtraction. The core idea extends to the case with value-function subtraction, except that separate value functions must be learned for the multi-fidelity environments because of the misspecified reward.
@@
These results support our key insight I3. Taken together, our experiments validate all three key insights (I1I3).
Figure 6: Evaluation return curves of MFPG, High-Fidelity Only, and Low-Fidelity Only on Hopper-v3, where the low-fidelity environment is misspecified with a negated reward. Solid lines indicate bootstrap point estimates (means), and shaded regions denote two-sided 95% bootstrap confidence intervals. Although low-fidelity data alone is ineffective for training, MFPG successfully leverages both fidelities to achieve substantial improvements over High-Fidelity Only in high-fidelity data-scarce regimes.
[…]
@@ 6.5 Additional Experiments
This section presents additional experiments that provide a more detailed examination of different aspects of the MFPG framework. For brevity in the main text, Appendix˜D presents additional ablation results.
MFPG performance with different amounts of high-fidelity samples.
@@
Our main evaluation in Section˜6.3 focuses on the high-fidelity data-scarce regime, which reflects many real-world scenarios where collecting large quantities of high-fidelity samples for each policy update is prohibitively expensive (e.g., in robotics or molecule design). This section studies how MFPG compares to the High-Fidelity Only baseline as the number of high-fidelity samples per policy update varies.
[…]
Figure 8: Mean difference in final return (top) and AUC (bottom) of MFPG with increasing amounts of low-fidelity data (light to dark orange) relative to High-Fidelity Only on Hopper-v3 with 0.5× and 1.2× friction (0 corresponds to High-Fidelity Only). Error bars indicate two-sided 95% bootstrap confidence intervals for the mean differences. MFPG is relatively robust to the amount of low-fidelity data: across the tested range of low-fidelity sample sizes, we do not observe statistically significant degradation relative to High-Fidelity Only under our significance test, and MFPG is often statistically significantly better.
Figure 9: Mean difference in final return (top) and AUC (bottom) of MFPG with varying exponential moving-average weights ($\eta_{\text{ma}}$ in Eq.˜3) relative to High-Fidelity Only (0 corresponds to High-Fidelity Only) on Hopper-v3 with 0.5× and 1.2× friction. Error bars indicate two-sided 95% bootstrap confidence intervals for the mean differences. Across the tested values of $\eta_{\text{ma}}$, MFPG is often statistically significantly better than High-Fidelity Only, and we do not observe statistically significant degradation.
MFPG performance with different amounts of low-fidelity samples.
@@
We assume throughout this work that low-fidelity samples are cheap to collect, and all methods are given access to an abundant budget of low-fidelity data. Consequently, optimizing the efficiency of low-fidelity sample usage (Cutler et al., 2015) is not the focus of this work. Instead, this section qualitatively examines how the performance of MFPG varies with different amounts of low-fidelity data.
@@
In principle, the performance of MFPG is robust with respect to the amount of low-fidelity samples, so long as the low-fidelity sample amount is sufficient to estimate the low-fidelity sample mean $\hat{\mu}^{l}$. Figure˜9 reports mean differences in final return (top) and AUC (bottom) of MFPG relative to High-Fidelity Only (0 corresponds to High-Fidelity Only) with increasing amounts of low-fidelity samples (indicated from light to dark orange) for estimating low-fidelity sample mean ($\hat{\mu}^{l}$ in Fig.˜1), in Hopper with 0.5× and 1.2× friction. Error bars indicate two-sided 95% bootstrap confidence intervals.
[…]
Sensitivity to exponential moving-average weight.
@@
We further examine the sensitivity of MFPG to the choice of the exponential moving-average weight $\eta_{\text{ma}}$ in Eq.˜3. Figure˜9 reports the mean differences in final return and AUC between MFPG (with different $\eta_{\text{ma}}$ values) and High-Fidelity Only (0 corresponds to High-Fidelity Only) in the Hopper task under 0.5× and 1.2× friction. Error bars indicate two-sided 95% bootstrap confidence intervals for the mean differences. Overall, across the tested values of $\eta_{\text{ma}}$, MFPG is often statistically significantly better than High-Fidelity Only, and we do not observe statistically significant degradation. The mean performance improvement of MFPG over High-Fidelity Only is generally more substantial in settings with smaller dynamics gaps—consistent with the key insights from our baseline comparison in Section˜6.3.